HUBBLE SPACE TELESCOPE DAILY REPORT #1540 PERIOD COVERED: 0000Z (UTC) 12/19/95 - 0000Z (UTC) 12/20/95 Daily Status Report as of 354/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 6339 (Verification of the UV Spectrum of the HST = Background Light ) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet observations of the HST background light. These FOS = observations were taken parallel to the Hubble Deep Field (HDF) survey. = Both the HDF and these FOS data will be used to test and improve the = existing engineering model of the HST background light. The observations = were completed as planned, and no anomalies were reported. 1.2 Completed WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration = images. Both photometric and narrow-band filters were used. The = observations were executed as planned, and no problems were reported. 1.3 Completed WF/PC-2 6337 (The Hubble Deep Field Observations) The WF/PC-2 were used to take the Hubble Deep Field (HDF) images. = This project consists of hundreds of long WF/PC-2 exposures of the high = galactic latitude field (HDF-124649), taken in eight visits. A dark = calibration image was also taken. The observations were completed as = planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 0 Successful : 0 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 3 Successful : 3 2.3 Operations Notes: There were no significant operational problems. HST operations transferred to the SAMS backup facility at = 354/1015Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH